you are provided KMnO4 solution which is about N/100. however you prepared standard solution of mohr salt which is about N/20, can 10ml of this standard mohr salt solution be used fot tiitrating with given KMnO4
1
Expert's answer
2020-12-21T03:57:21-0500

The overall chemical equation in titration of Mohr salt with KMnO4 is

"2KMnO_4 + 10FeSO_4(NH_4)_2SO_4.6H_2O+8H_2O \to K_2SO_4+ 2MnSO_4+ 5Fe_2(SO_4)_3+ 10(NH_4)_2SO_4+ 68H_2O"


number of moles of Mohr Salt used = normality × volume = 1/20 × 10/1000 = 1/2000 moles.


from the reaction above, 10 moles of Mohr Salt reacts with 2 moles of KMnO4.

Therefore, 1/2000 moles of Mohr Salt reacts with 1/10000 moles of KMnO4.


volume of KMnO4 = number of moles/normality

= 1/10000 ÷ 1/100 = 1/100L = 10mL


This means that 10ml of this standard mohr salt solution be used fot tiitrating with given 10ml of this standard mohr salt solution be used fot tiitrating with given KMnO4 solution, provided it's volume is 10mL or above.